steve is three times as old as bubut. in 8 years, bubuts age will be one half that of steve's. what are their present ages?
Answer by mananth(16946)

bubut = x years
steve = 3x
...
after 8 years
bubut = x+ 8 years
steve = 3x+8
x+8 = 1/2(3x+8)
2(x+8)=3x+8
2x+16=3x+8
3x-2x=16-8
x= 8 bubut's age
steve will be 3x = 24 years
..
m.ananth@hotmail.ca